I have a rectangle with an area of 200cm2
I make the length 20% longer and the width 20% shorter
What is the new area?

Answer by Alan3354(69443)

I have a rectangle with an area of 200cm2
I make the length 20% longer and the width 20% shorter
What is the new area?
-------
Area = L*W = 200
New area = 0.8L*1.2W = 0.96*200 = 192